# Black-Box Access is Insufficient for Rigorous AI Audits

### v1

@@ Abstract.
External audits of AI systems are increasingly recognized as a key mechanism for AI governance. The effectiveness of an audit, however, depends on the degree of system access granted to auditors. Recent audits of state-of-the-art AI systems have primarily relied on black-box access, in which auditors can only query the system and observe its outputs. However, white-box access to the system’s inner workings (e.g., weights, activations, gradients) allows an auditor to perform stronger attacks, more thoroughly interpret models, and conduct fine-tuning. Meanwhile, outside-the-box access to its training and deployment information (e.g., methodology, code, documentation, hyperparameters, data, deployment details, findings from internal evaluations) allows for auditors to scrutinize the development process and design more targeted evaluations. In this paper, we examine the limitations of black-box audits and the advantages of white- and outside-the-box audits. We also discuss technical, physical, and legal safeguards for performing these audits with minimal security risks. Given that different forms of access can lead to very different levels of evaluation, we conclude that (1) transparency regarding the access and methods used by auditors is necessary to properly interpret audit results, and (2) white- and outside-the-box access allow for substantially more scrutiny than black-box access alone.
@@
copyright: acmcopyrightjournalyear: 2024ccs: Security and privacy Social aspects of security and privacyccs: Social and professional topics Governmental regulations
@@ 1. Introduction
External evaluations of AI systems are emerging as a key component of AI oversight (Brown et al., 2021; Watkins et al., 2021; Metcalf et al., 2021; Liu et al., 2022a; Metcalf et al., 2022; Raji et al., 2022; Raji, 2022; Raji and Buolamwini, 2022; Koshiyama et al., 2022; Schuett, 2022; Schuett et al., 2023; Mökander et al., 2023a; Shevlane et al., 2023; Seger et al., 2023; Landers and Behrend, 2023; Anderljung et al., 2023a; Solaiman, 2023; Solaiman et al., 2023; Wilkinson et al., 2023; Mökander, 2023; Miotti and Wasil, 2023; Lee et al., 2023; Anderljung et al., 2023b; Bengio et al., ; METR, 2023) and governance frameworks (European Commission, 2021; U.S. Department of Commerce and National Institute of Standards and Technology, 2023; Chinese National Information Security Standardization Technical Committee, 2023; UK Department for Science, Innovation & Technology, 2023; Office of the President of the United States, 2023). There is a rich history in academic AI research of evaluating systems to explain their behaviors (Zhang et al., 2022; Agarwal et al., 2022) and evaluate risks including those related to privacy (Yao et al., 2023; Smith et al., 2023; Carlini et al., 2021; Singhal et al., 2022; Shi et al., 2023), intellectual property rights (Karamolegkou et al., 2023; Casper et al., 2023a; Carlini et al., 2022), fairness and discrimination (Bolukbasi et al., 2016; Buolamwini and Gebru, 2018; Eckhouse et al., 2019; Coe and Atay, 2021; Mehrabi et al., 2021; Weidinger et al., 2021; Angwin et al., 2022; Gichoya et al., 2022; Chen et al., 2022; Tao et al., 2023; Dhamala et al., 2021; Krishna et al., 2022), harmful content, (Birhane et al., 2021, 2023; Thiel, 2023; Thiel et al., 2023; Qu et al., 2023; Rando et al., 2022), circumvention of safeguards (“jailbreaks”) (Shayegani et al., 2023; Liu et al., 2023a; Rao et al., 2023; Wei et al., 2023; Zou et al., 2023a; Shah et al., 2023; Carlini et al., 2023a), misinformation and deception (Ji et al., 2023; Huang et al., 2023; Scheurer et al., 2023; Park et al., 2023a; Hubinger et al., 2024), dangerous capabilities (Charan et al., 2023; Shevlane et al., 2023; Kinniment et al., 2023; Chan et al., 2023; OpenAI, 2023a), and broader societal impacts (Solaiman et al., 2023; Weidinger et al., 2023).
@@
Historically, most academic work on evaluating AI systems has been conducted on models where parameters, data, and methodology are openly available. AI systems that are not available to the public, including ones that are proprietary or in pre-deployment, pose challenges for oversight. AI audits are structured evaluations designed to identify risks and improve transparency by assessing how well models and methods meet specific desiderata (Costanza-Chock et al., 2022; Mökander et al., 2023b). Norms for AI audits are not yet well established, and their effectiveness can vary depending on the degree of system access granted to auditors (Brundage et al., 2020; Raji et al., 2022; Anderljung et al., 2023b). This is crucial because existing calls for audits are often agnostic to the form of access, and industry actors have previously lobbied for limiting access given to auditors (Google, 2021).
@@
Recently, some developers of prominent state-of-the-art AI systems have kept most details of their models private (Bommasani et al., 2023). To public knowledge, voluntary external audits of these systems have primarily involved analysis of the input/output behavior of models (METR, 2023; OpenAI, 2023b; Anthropic, 2023; Touvron et al., 2023). This form of access, in which auditors are only able to see outputs for given inputs, is known as black-box. Unfortunately, black-box access is very limiting for auditors. Some problems, such as anomalous failures, are difficult to find with black-box access (Kolt, 2023), and others, such as dataset biases, can be actively reinforced by testing data (Shahbazi et al., 2023).
@@
The ability to query a black-box system is useful, but many of today’s evaluation techniques require access to weights, activations, gradients, or the ability to fine-tune the model (Bucknall and Trager, 2023). White-box access refers to the unrestricted ability to observe a system’s internal workings. It enables evaluators to apply more powerful attacks to automatically identify weaknesses (Goodfellow et al., 2014; Papernot et al., 2016), study internal mechanisms responsible for undesirable model behaviors (Jain et al., 2023; Lee et al., 2024), and identify harmful dormant capabilities through fine-tuning (Qi et al., 2023a; Zhan et al., 2023). Meanwhile, outside-the-box access involves additional contextual information about a system’s development or deployment such as methodology, code, documentation, hyperparameters, data, deployment details, and findings from internal evaluations. It allows auditors to study risks that stem from methodology or data (Mitchell et al., 2019; Birhane et al., 2021; Shahbazi et al., 2023; Casper et al., 2023b) and makes it easier to design useful tests. This paper makes four contributions:
@@
Given the growing evidence that different forms of access can facilitate very different levels of evaluation, we draw two conclusions. First, transparency regarding model access and evaluation methods is necessary to properly interpret the results of an AI audit. Second, white- and outside-the-box access allow for substantially more scrutiny than black-box access alone. When higher levels of scrutiny are desired, audits should be conducted with higher levels of access.
@@ 2.1. Black, Grey, White, and Outside-the-Box Access
Identifying problems Attacks
White-box methods are more reliable for detecting anomalous failures.
@@
Fine-tuning
Fine-tuning enables searching for harmful dormant capabilities.
Interpretability
Interpretability tools enable stronger assurances of what knowledge and capabilities a system has.
Outside-the-box assessment
Information helps auditors target complementary evaluations.
Incentivizing responsible development Outside-the-box assessment
Information about deployment assists with assessing societal impacts.
@@
Increasing transparency Interpretability
White-box methods can address misconceptions about model outputs, characteristics, and risks.
@@
Enabling debugging Attacks
Stronger attacks produce stronger instances of failures to incorporate into training data.
Interpretability
Interpretability techniques enable more precise debugging.
@@ 2.2. Regulatory Frameworks’ Reliance on Audits
Assessing model explanations enables scrutiny of automated decisions. Regulatory frameworks also use audits to provide those affected by automated decision-making with explanations of the decisions (Gryz and Rojszczak, 2021; Ploug and Holm, 2021). In some jurisdictions, such as the European Union, when individuals are harmed by automated decision-making systems, they may have the right to an explanation, and the results of these explanations may entitle them to remediation (Hacker and Passoth, 2022; Buiten et al., 2023). Further, explanation requirements may exist for particularly high-risk systems, such as EU platform regulations (the Digital Markets Act and the Digital Services Act) that require transparency from large online platforms using AI systems (e.g., ranking algorithms) to protect against discrimination and abuse of market power (European Union, 2022; Hacker et al., 2023). Finally, disclosure of evidence may be required under liability rules, such as the Product and AI Liability Directives in the EU, to enable potential claimants to adequately defend damage claims (Buiten et al., 2021; Hacker, 2023). When producing explanations, a report from NIST emphasizes the importance of explanation accuracy, or “[a]n explanation [that] correctly reflects the reason for generating the output and/or accurately reflects the system’s process” (Phillips et al., 2021b). In Section 3 and Section 4.2, we overview advantages of white-box access over black-box access for generating reliable explanations.
@@ 2.3. Audits in the Status Quo
Recent advancements in AI capabilities – especially from large generative models – have increased public attention on AI audits. As of January 2024, there are no widely adopted norms for conducting AI audits. Details of AI audits can vary because they depend upon the system, how it will be used, and what risks it poses. Auditing frameworks and metrics have been proposed for specific use cases, including hiring (Kazim et al., 2021; Raghavan et al., 2020), facial recognition (Khalil et al., 2020; Raji et al., 2020a), healthcare (Mahajan et al., 2020; Liu et al., 2022a), recommender systems (Robertson et al., 2018; Chen et al., 2023), and general purpose language models (Mökander et al., 2023b). However, Raji et al. (2022) identifies five general limitations for algorithmic audits: scope, independence, level of access, professionalism, and public disclosure of methods and results.
Currently, evaluations of proprietary or pre-deployment AI systems are predominantly performed in-house by developers with selective disclosure of methods and outcomes. Some developers have voluntarily partnered with external auditors and provided them with black-box access to state-of-the-art systems (METR, 2023; OpenAI, 2023a; Anthropic, 2023; Touvron et al., 2023; Kinniment et al., 2023). Additionally, some developers run programs for external researchers to support their internal evaluation process (e.g., OpenAI’s Preparedness Challenge (OpenAI, 2023c) and Red-Teaming Network (OpenAI, 2023d)). However, to public knowledge, these industry-based efforts have involved black-box and limited outside-the-box access such as “model cards”, (Mitchell et al., 2019).
@@ 3. Limitations of Black-Box Access
Black-box evaluations of AI systems are based on analysis of their inputs and outputs only. Such evaluations often involve assessing performance on test sets (Liang et al., 2022; Sarlin et al., 2020; Wang et al., 2018; Srivastava et al., 2022; Hendrycks et al., 2020; Sun et al., 2024) or searching for inputs that elicit harmful outputs (Perez et al., 2022a, b; Casper et al., 2023c; Kinniment et al., 2023; Scheurer et al., 2023; Wei et al., 2023). Generative AI audits often attempt to elicit undesirable capabilities or behaviors (e.g., (Mouton et al., 2023; Sandbrink, 2023; Tann et al., 2023; Bran et al., 2023; Kinniment et al., 2023; Soice et al., 2023; Shevlane et al., 2023; Hazell, 2023; Karanjai, 2022; Nasr et al., 2023)). However, black-box methods are inherently limited in their ability to identify harms or provide meaningful explanations. Readers with a computer science background can consider the analogy of attempting to evaluate the performance of software without reading or modifying its source code.
@@
Black-box access prevents system components from being studied separately. Analyzing components of a system separately is ubiquitous in science and engineering. It enables engineers to trace problems to support more targeted interventions. However, black-box access obscures what subsystems the AI system is composed of. For example, black-box access does not allow input or output filters to be studied separately from the rest of the system. Other issues can arise from a lack of outside-the-box access to data. Datasets can inform evaluations related to privacy and copyright (Roth, 2023), and can help to avoid problems from data contamination (Deng et al., 2023a; Jacovi et al., 2023; Golchin and Surdeanu, 2023). Having a lack of outside-the-box knowledge about how the system is deployed also prevents evaluators from making a more practical assessment of broader societal impacts (Weidinger et al., 2023).
Black-box evaluations cand produce misleading results. Since black-box evaluations rely entirely on the queries made to the system, they are biased by how evaluators design inputs (Wei et al., 2022; Kapoor and Narayanan, 2023; Narayanan and Kapoor, 2023). This can lead to misconstrued conclusions about the system’s characteristics. For example, systems may satisfy simple statistical tests for non-discrimination, but they may still have undesirable biases in their underlying reasoning (Raghavan and Kim, 2023). In addition, Schaeffer et al. (2023) provide examples of black-box prompt-based evaluation methods for language models that can lead to misunderstandings of their emergent capabilities.
@@
Black-box evaluations offer limited insights to help address failures. Black-box evaluations offer little insight into ways to address problems they discover. The main technique they enable is to train on problematic examples, but this can fail to address the underlying problem (Geirhos et al., 2020a; Du et al., 2023), be sample-inefficient (Wang et al., 2023a), and may introduce new issues. Corrective actions are not robust when they fail to address a problem at its root. For example, some recent works have shown that safety measures built into large language models can be almost entirely undone by fine-tuning on a small number of harmful examples (Yang et al., 2023; Qi et al., 2023a; Lermen et al., 2023; Zhan et al., 2023). In contrast, white-box methods reveal more about the nature of flaws, facilitating more precise debugging methods (Wang et al., 2023a).

### v2

@@ Abstract.
External audits of AI systems are increasingly recognized as a key mechanism for AI governance. The effectiveness of an audit, however, depends on the degree of access granted to auditors. Recent audits of state-of-the-art AI systems have primarily relied on black-box access, in which auditors can only query the system and observe its outputs. However, white-box access to the system’s inner workings (e.g., weights, activations, gradients) allows an auditor to perform stronger attacks, more thoroughly interpret models, and conduct fine-tuning. Meanwhile, outside-the-box access to training and deployment information (e.g., methodology, code, documentation, data, deployment details, findings from internal evaluations) allows auditors to scrutinize the development process and design more targeted evaluations. In this paper, we examine the limitations of black-box audits and the advantages of white- and outside-the-box audits. We also discuss technical, physical, and legal safeguards for performing these audits with minimal security risks. Given that different forms of access can lead to very different levels of evaluation, we conclude that (1) transparency regarding the access and methods used by auditors is necessary to properly interpret audit results, and (2) white- and outside-the-box access allow for substantially more scrutiny than black-box access alone.
@@
journalyear: 2024copyright: acmcopyrightconference: The 2024 ACM Conference on Fairness, Accountability, and Transparency; June 3–6, 2024; Rio de Janeiro, Brazilbooktitle: The 2024 ACM Conference on Fairness, Accountability, and Transparency (FAccT ’24), June 3–6, 2024, Rio de Janeiro, Brazildoi: 10.1145/3630106.3659037isbn: 979-8-4007-0450-5/24/06doi: 3630106.3659037conference: ; June 03–06, 2024; Rio de Janeiro, Brazilccs: Security and privacy Social aspects of security and privacyccs: Social and professional topics Governmental regulations
@@ 1. Introduction
External evaluations of AI systems are emerging as a key component of AI oversight (Brown et al., 2021; Watkins et al., 2021; Metcalf et al., 2021; Liu et al., 2022a; Metcalf et al., 2022; Raji et al., 2022; Raji, 2022; Raji and Buolamwini, 2022; Koshiyama et al., 2022; Schuett, 2022; Schuett et al., 2023; Mökander et al., 2023a; Shevlane et al., 2023; Seger et al., 2023; Landers and Behrend, 2023; Anderljung et al., 2023a; Solaiman, 2023; Solaiman et al., 2023; Wilkinson et al., 2023; Mökander, 2023; Miotti and Wasil, 2023; Lee et al., 2023; Anderljung et al., 2023b; Bengio et al., [n. d.]; Sharkey et al., 2024; Birhane et al., 2024; METR, 2023) and governance frameworks (European Commission, 2021; U.S. Department of Commerce and National Institute of Standards and Technology, 2023; Chinese National Information Security Standardization Technical Committee, 2023; UK Department for Science, Innovation & Technology, 2023; Office of the President of the United States, 2023). There is a rich history in academic AI research of evaluating systems to explain their behaviors (Zhang et al., 2022; Agarwal et al., 2022) and evaluate risks including those related to privacy (Yao et al., 2023; Smith et al., 2023; Carlini et al., 2021; Singhal et al., 2022; Shi et al., 2023), intellectual property rights (Karamolegkou et al., 2023; Casper et al., 2023b; Carlini et al., 2022), fairness and discrimination (Bolukbasi et al., 2016; Buolamwini and Gebru, 2018; Eckhouse et al., 2019; Coe and Atay, 2021; Mehrabi et al., 2021; Weidinger et al., 2021; Angwin et al., 2022; Gichoya et al., 2022; Chen et al., 2022; Tao et al., 2023; Dhamala et al., 2021; Krishna et al., 2022), harmful content, (Birhane et al., 2021, 2023; Thiel, 2023; Thiel et al., 2023; Qu et al., 2023; Rando et al., 2022), circumvention of safeguards (jailbreaks) (Shayegani et al., 2023; Liu et al., 2023a; Rao et al., 2023; Wei et al., 2023; Zou et al., 2023b; Shah et al., 2023; Carlini et al., 2023b), misinformation and deception (Ji et al., 2023; Huang et al., 2023; Scheurer et al., 2023; Park et al., 2023a; Hubinger et al., 2024), dangerous capabilities (Charan et al., 2023; Shevlane et al., 2023; Kinniment et al., 2023; Chan et al., 2023; OpenAI, 2023b), and broader societal impacts (Solaiman et al., 2023; Weidinger et al., 2023).
@@
Historically, most academic work on evaluating AI systems has been conducted on models where parameters, data, and methodology are openly available. AI systems that are not available to the public, including ones that are proprietary or in pre-deployment, pose challenges for oversight. AI audits are structured evaluations designed to identify risks and improve transparency by assessing how well models and methods meet specific desiderata (Costanza-Chock et al., 2022; Mökander et al., 2023b). Norms for AI audits are not yet well established, and their effectiveness can vary depending on the degree of system access granted to auditors (Brundage et al., 2020; Raji et al., 2022; Anderljung et al., 2023b). This is crucial because existing calls for audits are often agnostic to the form of access, and industry actors have previously lobbied for limiting access given to auditors. [^1] For example, Google made the unsubstantiated claim in 2021 that “There will always be better methods for verifying the performance of an AI system (ex: input/output auditing) than direct access to source code,” while lobbying the European Union on a draft of the EU AI Act (Google, 2021).
@@
Recently, some developers of prominent state-of-the-art AI systems have kept most details of their models private (Bommasani et al., 2023). To public knowledge, voluntary external audits of these systems have primarily involved analysis of the input/output behavior of models (METR, 2023; OpenAI, 2023a; Anthropic, 2023; Touvron et al., 2023). This form of access, in which auditors are only able to see outputs for given inputs, is known as black-box. Unfortunately, black-box access is very limiting for auditors. Some problems, such as anomalous failures, are difficult to find with black-box access (Kolt, 2023), and others, such as dataset biases, can be actively reinforced by testing data (Shahbazi et al., 2023).
@@
The ability to query a black-box system is useful, but many of today’s evaluation techniques require access to weights, activations, gradients, or the ability to fine-tune the model (Bucknall and Trager, 2023). White-box access refers to the unrestricted ability to observe a system’s internal workings. It enables evaluators to apply more powerful attacks to automatically identify weaknesses (Goodfellow et al., 2014; Papernot et al., 2016), study internal mechanisms responsible for undesirable model behaviors (Jain et al., 2023; Lee et al., 2024; Krishna et al., 2024), and identify harmful dormant capabilities through fine-tuning (Qi et al., 2023b; Zhan et al., 2023). Meanwhile, outside-the-box access involves additional contextual information about a system’s development or deployment such as methodology, code, documentation, hyperparameters, data, deployment details, and findings from internal evaluations. It allows auditors to study risks that stem from methodology or data (Mitchell et al., 2019; Birhane et al., 2021; Shahbazi et al., 2023; Casper et al., 2023a) and makes it easier to design useful tests. This paper makes four contributions:
@@
Given the growing evidence that different forms of access can facilitate very different levels of evaluation, we draw two conclusions. First, transparency regarding model access and evaluation methods is necessary to properly interpret the results of an AI audit. Second, white- and outside-the-box access allow for substantially more scrutiny than black-box access alone. When higher levels of scrutiny are desired, audits should be conducted with higher levels of access. [^2] Jointly with this paper, we submitted a comment on policy questions related to white- and outside-the-box audits to a request for information from the US National Institute of Standards and Technology (National Institute for Standards and Technology, 2023).
@@ 2.1. Black, Grey, White, and Outside-the-Box Access
Identifying problems Attacks White-box methods are more reliable for detecting anomalous failures.
@@
Fine-tuning Fine-tuning enables searching for harmful dormant capabilities.
Interpretability Interpretability tools enable stronger assurances of what knowledge and capabilities a system has.
Outside-the-box assessment Information helps auditors target complementary evaluations.
Incentivizing responsible development Outside-the-box assessment Information about deployment assists with assessing societal impacts.
@@
Increasing transparency Interpretability White-box methods can address misconceptions about model outputs, characteristics, and risks.
@@
Enabling debugging Attacks Stronger attacks produce stronger instances of failures to incorporate into training data.
Interpretability Interpretability techniques enable more precise debugging.
@@ 2.2. Regulatory Frameworks’ Reliance on Audits
Assessing model explanations enables scrutiny of automated decisions. Regulatory frameworks also use audits to provide those affected by automated decision-making with explanations of the decisions (Gryz and Rojszczak, 2021; Ploug and Holm, 2021). In some jurisdictions, such as the European Union, when individuals are harmed by automated decision-making systems, they may have the right to an explanation, and the results of these explanations may entitle them to remediation (Hacker and Passoth, 2022; Buiten et al., 2023). Further, explanation requirements may exist for particularly high-risk systems, such as EU platform regulations (the Digital Markets Act and the Digital Services Act) that require transparency from large online platforms using AI systems (e.g., ranking algorithms) to protect against discrimination and abuse of market power (European Union, 2022; Hacker et al., 2023). Finally, disclosure of evidence may be required under liability rules, such as the Product and AI Liability Directives in the EU, to enable potential claimants to adequately defend damage claims (Buiten et al., 2021; Hacker, 2023). When producing explanations, a report from NIST emphasizes the importance of explanation accuracy, or [a]n explanation [that] correctly reflects the reason for generating the output and/or accurately reflects the system’s process (Phillips et al., 2021b). In Section 3 and Section 4.2, we overview advantages of white-box access over black-box access for generating reliable explanations.
@@ 2.3. Audits in the Status Quo
Recent advancements in AI capabilities – especially from large generative models – have increased public attention on AI audits. As of January 2024, there are no widely adopted norms for conducting AI audits (Birhane et al., 2024). Details of AI audits can vary because they depend upon the system, how it will be used, and what risks it poses. Auditing frameworks and metrics have been proposed for specific use cases, including hiring (Kazim et al., 2021; Raghavan et al., 2020), facial recognition (Khalil et al., 2020; Raji et al., 2020a), healthcare (Mahajan et al., 2020; Liu et al., 2022a), recommender systems (Robertson et al., 2018; Chen et al., 2023), and general purpose language models (Mökander et al., 2023b). However, Raji et al. (2022) identifies five general limitations for algorithmic audits: scope, independence, level of access, professionalism, and public disclosure of methods and results.
Currently, evaluations of proprietary or pre-deployment AI systems are predominantly performed in-house by developers with selective disclosure of methods and outcomes. Some developers have voluntarily partnered with external auditors and provided them with black-box access to state-of-the-art systems (METR, 2023; OpenAI, 2023b; Anthropic, 2023; Touvron et al., 2023; Kinniment et al., 2023). Additionally, some developers run programs for external researchers to support their internal evaluation process (e.g., OpenAI’s Preparedness Challenge (OpenAI, 2023c) and Red-Teaming Network (OpenAI, 2023d)). However, to public knowledge, these industry-based efforts have involved black-box and limited outside-the-box access such as model cards (Mitchell et al., 2019).
@@ 3. Limitations of Black-Box Access
Black-box evaluations of AI systems are based on analysis of their inputs and outputs only. Such evaluations often involve assessing performance on test sets (Liang et al., 2022; Sarlin et al., 2020; Wang et al., 2018; Srivastava et al., 2022; Hendrycks et al., 2020; Sun et al., 2024) or searching for inputs that elicit harmful outputs (Perez et al., 2022b, a; Casper et al., 2023e; Kinniment et al., 2023; Scheurer et al., 2023; Wei et al., 2023). Generative AI audits often attempt to elicit undesirable capabilities or behaviors (e.g., (Mouton et al., 2023; Sandbrink, 2023; Tann et al., 2023; Bran et al., 2023; Kinniment et al., 2023; Soice et al., 2023; Shevlane et al., 2023; Hazell, 2023; Karanjai, 2022; Nasr et al., 2023)). However, black-box methods are inherently limited in their ability to identify harms or provide meaningful explanations. Readers with a computer science background can consider the analogy of attempting to evaluate the performance of software without reading or modifying its source code.
@@
Black-box access prevents system components from being studied separately. Analyzing components of a system separately is ubiquitous in science and engineering. It enables engineers to trace problems to support more targeted interventions. However, black-box access obscures what subsystems the AI system is composed of. For example, black-box access does not allow input or output filters to be studied separately from the rest of the system. Other issues can arise from a lack of outside-the-box access to data. Datasets can inform evaluations related to privacy and copyright (Roth, 2023), and can help to avoid problems from data contamination (Deng et al., 2023b; Jacovi et al., 2023; Golchin and Surdeanu, 2023). Having a lack of outside-the-box knowledge about how the system is deployed also prevents evaluators from making a more practical assessment of broader societal impacts (Weidinger et al., 2023).
Black-box evaluations can produce misleading results. Since black-box evaluations rely entirely on the queries made to the system, they are biased by how evaluators design inputs (Wei et al., 2022; Kapoor and Narayanan, 2023; Narayanan and Kapoor, 2023). This can lead to misconstrued conclusions about the system’s characteristics. For example, systems may satisfy simple statistical tests for non-discrimination, but they may still have undesirable biases in their underlying reasoning (Raghavan and Kim, 2023). Developers with information about the black-box tests can exacerbate this problem by modifying the model’s output behavior on test cases despite unresolved flaws in its internal reasoning. In addition, Schaeffer et al. (2023) provide examples of black-box prompt-based evaluation methods for language models that can lead to misunderstandings of their emergent capabilities.
@@
Black-box evaluations offer limited insights to help address failures. Black-box evaluations offer little insight into ways to address problems they discover. The main technique they enable is to train on problematic examples, but this can fail to address the underlying problem (Geirhos et al., 2020a; Du et al., 2023), be sample-inefficient (Wang et al., 2023c), and may introduce new issues. Corrective actions are not robust when they fail to address a problem at its root. For example, some recent works have shown that safety measures built into large language models can be almost entirely undone by fine-tuning on a small number of harmful examples (Yang et al., 2023; Qi et al., 2023b; Lermen et al., 2023; Zhan et al., 2023). In contrast, white-box methods reveal more about the nature of flaws, facilitating more precise debugging methods (Wang et al., 2023c).